A car accelerates from rest at a constant rate of 2.0 m/s for 5.0 s. (a) What is the speed of the car at the end of that time? (b) How far does the car travel in this time? MacBook Pro
Answer #1

Given things:

u = 0 m/s (initial speed)

a = 2 m/s^2

t = 5 sec

A)

by using the kinematical equation

v = u + at v is the velocity after time t

v = 0 + 2 x 5 =

v = 10 m/s

B)

by using the third kinematical equation

s = ut + 0.5 at2

where s = displacement

s = 0 + 0.5 x 2 x 52

s = 25 m